A bottle full of meecury has a mass of 720gms.When full of water its mass is 90gms. Calculate the volume of empty density bottle.


1
Expert's answer
2021-10-10T16:01:28-0400

The mass of the bottle full of mercury and water:


"M_1=m+\rho_1V,\\
M_2=m+\rho_2V."

Subtract one from another:


"M_1-M_2=V(\rho_1-\rho_2),\\\space\\
V=\frac{M_1-M_2}{\rho_1-\rho_2}=5·10^{-5}\text{ m}^3,"

or 0.05 L.

Mass of the bottle:


"m=M_1-\rho_1V=0.04\text{ kg}."


Density of the bottle:


"\rho=\frac{m}{V}=800\text{ kg/m}^3."